The sum of two numbers is 22 . The sum of 6 times the larger and 7 times the smaller is 137. What are x and y?
Answer by ikleyn(52787)

.
The sum of two numbers is 22 . The sum of 6 times the larger and 7 times the smaller is 137.
What are x and y?
~~~~~~~~~~~~~~~


            Notice that both terms  "x"  and  "y"  are not defined in your post.

            It is a  HUGE  DEFICIENCY  of your creation.


Let x be the larger number; then the smaller number is  (22-x).


Write the equation for the sum of 137 as you read the post


    6x + 7*(22-x) = 137.


Simplify and solve


    6x + 152 - 7x = 137

    -x            = 137 - 152

    -x            = -17

     x            = 17.


ANSWER.  The numbers are  17 and 22-17 = 5.


CHECK.   6*17 + 7*5 = 137.

Solved, checked and explained.